OpenStudy (anonymous):

Solve: 4(x + 6) = 6(x − 2) Solve: 2(x + 1) = −4 Solve: 3x = −18 Solve: 2 − x = 13 Solve: 2(x + 4) − (x + 1) = 10 Solve: x − 5 = 10 lovee you guys /.\ lol

OpenStudy (anonymous):

ill work these out with u :) wat did u get for the first question?

OpenStudy (anonymous):

so im supposed to add 6 or something right? cause i havee no idea what to do yet lol..

OpenStudy (anonymous):

ok first expand the brackets :)

OpenStudy (anonymous):

and u dont add, when expanding brackets u multiply wat's outside of the brackets with wat is inside of the brackets, like this 4(x + 6) = 6(x − 2) 4x + 24 = 6x -12

OpenStudy (anonymous):

now solve for x given that :)

OpenStudy (ryan123345):